Science Report – April 16th

This afternoon crew 210 began to conduct vegetation quadrat surveys as a part of our local plant and lichen biodiversity survey. At two sites pre-selected from the MDRS topographic map (both east of the hab along Cactus Road) we plotted three 10-meter transects, and surveyed five 1-meter-square quadrats along each transect in a zig-zag pattern.

For each transect we recorded GPS coordinates, air and soil
temperature at the center of the transect, measured the height of the tallest plant on the transect, and took a central sample for soil moisture and pH (these latter variables will be measured in the lab tomorrow).

Within each of the five quadrats, we estimated percent cover of vascular plants, and named the top three species (by abundance) inside the square-meter. Any new plant species encountered were collected for identification and to add to our plant biodiversity data.

Tomorrow we will finish the quadrat surveys at three additional pre-determined sites, north and west of the hab, and will continue our plant and lichen collection efforts.

Paul, on behalf of Crew 210

Greenhab Report – April 16th

GreenHab Officer: David Murray

Environmental control: Ambient with door open 9:35 to 17:20

Shade Cloth: 80%

Low Temperature: 16 C

High Temperature: 26 C

Relative humidity: 22%

Hours of Supplemental Light: 0

Daily Water Usage for Crops: 32 L

Water in Blue Tank: Low, needs refilling

Time(s) of Watering: 9:40

Changes to Crops: None.

Narrative: All crops except for seedlings were fertilized with Miracle Grow today. It was much cooler in the GreenHab than it’s been the last few days. Did not use the fan. 40 g basil was harvested.

Harvest: 40 g basil

Sol Summary – April 16th

Crew 201 Sol Summary Report 16-APR-2019

Sol: 4

Summary Title: Quadrat by quadrat, we’re painting a picture of the “Martian” desert

Author’s name: Paul Sokoloff

Mission Status: Nominal.

Sol Activity Summary: (This morning we hosted a representative from the local Bureau of Land Management Office, therefore our team left for EVA in the early afternoon. We visited two sites east of the hab along the Cactus Road to survey vegetation using quadrats at two pre-determined sites – details for this sampling will be covered in the science report. At these sites we made additional plant and lichen collections for our biodiversity survey project. At the end of the day we proceeded to Hanksville to host an “Ask a Martian Anything” on Reddit.

Look Ahead Plan: Tomorrow we plan on visiting three more sites to finish our quadrat surveys, and to process the plants collected today in the hab science dome.

Anomalies in work: No issues to report.

Weather: Good weather (though slightly cold and overcast) until we returned home from our EVA, and which point rain and hail started over us.

Crew Physical Status: Nominal.

EVA: We drove along Cactus Road east of the station and sampled two sites along the route for our quadrat surveys.

Reports to be file: Operations Report, GreenHab Report, Photo Report, Science Report

Support Requested: None.

Operations Report – April 16th

Crew 210 Operations Report 16-APR-2019

SOL: 4

Name of person filing report: Shannon Rupert

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run): 10.5

Solar— SOC 98 % as of 5:30 pm. Will not turn on generator until needed.

Diesel Reading – 38 %

Propane Reading – 43 %

Ethanol Free Gasoline – 7 gallons.

Water (auxillary tank) – 0 gallons

Water (static tank) – 300 gallons

Auxillary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – unknown gallons.

Water (loft) – Static to Loft Pump used – yes

Water Meter: not read

Toilet tank emptied: no

Deimos rover used: Still non-operational


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, yes

Hours: Did not record

Beginning charge: 100

Ending charge: 49

Currently charging: yes

Spirit rover used: yes

Hours: did not check

Beginning charge: 100

Ending charge: 42

Currently charging: yes

Opportunity rover used: no

Hours: 56.4

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Curiosity rover used: No

Hours: 88.8

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) all of them

Reason for use: Testing and Field work

Oil Added? Yes. The oil in the 300 was fine. Oil in 350.1 and 350.3 was low but ok, but it was very low in 350.2 (the dirty one). The Honda had NO oil at all.

ATV Fuel Used: 3 Gals.

# Hours the ATVs were Used today: 5

Notes on ATVs: The battery was replaced in the 300 and it is now operational as well. We changed the spark plug in the Honda and it still wouldn’t start. Turned out the kill switch was on. When the kill switch in on in the Yamaha’s they won’t even turn over, in the Honda it will turn over but not engage the engine. At any rate, with no oil we would have ruined it if we had run it before we put oil in it, so all is good. The new spark plug wasn’t a waste, as it starts right up now.

When testing the ATVs, we discovered that the handlebars on 350.2 are now so low that it is damaging the machine, once again causing accelerator on the turn and reducing the turning ratio. We ran it into a rover due to these issues. The rover was undamaged but I decided to pull the ATV from use until I can figure out how to fix the
handlebars. If we do use this AYV, I will be the one to operate it/

HabCar used and why, where? No

CrewCar used and why, where? Yes, to Hanksville for AMA REDDIT thing.

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: We are having some issues with the wall plugs on the wall in the main space of the lower deck. This should be addressed in the near future.

Summary of GreenHab operations: David will send a full report today. He continues working on climate control.

Summary of ScienceDome operations: We collected many plants today and will be spending time in the dome later tonight and tomorrow processing them.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing tonight, thank you.

Copyright © The Mars Society. All rights reserved. | Main Site